# Semantic Trajectory Data Mining with LLM-Informed POI Classification

### v1

@@ IV-B2 Activity Inference
To assess the efficacy of our proposed activity inference model, we curate a test trajectory dataset by extracting activities from the 2017 National Household Travel Survey (NHTS) California Add-on dataset [37], focusing on LA County. This dataset provides detailed daily travel trajectories, with each trajectory consisting of activities and a rough location at the zonal level. However, due to the lack of POIs associated with these activities, we utilize LLM for preliminary POI selection from the rough location zone, followed by manual curation to rectify any erroneous assignments. It is important to note the limitations of GPS-enabled smartphones, which typically provide accurate location data within a 4.9-meter radius under ideal conditions [38]. However, accuracy diminishes in urban environments near structures like buildings, bridges, and dense foliage. To evaluate the model’s performance and robustness against real-world conditions, we introduce noise at three levels: 5 meters, 10 meters, and 20 meters. The noise, simulated using a Gaussian kernel, is added to the latitude and longitude coordinates of the POIs. Subsequent to this augmentation, our test dataset comprises 362 individuals and 2,007 activities, consisting of 1,724 mandatory activities and 283 non-mandatory activities.
@@ IV-C1 POI Classification
As shown in Table II, our model achieves a 93.4% accuracy for the Egypt dataset and 91.4% accuracy for the LA county dataset, with F-1 scores of 96.10% and 92.9% respectively when using the gpt4 model. This demonstrates its superiority and robustness compared with previous learning-based methods.
@@ IV-C2 Activity Inference
Figure 5: Semantic trajectories examples after data mining. (a) Activity-annotated stay points of trajectories in LA. (b) Activity timelines of trajectories.
According to Table III, which outlines the performance metrics for activity inference under varying levels of noise, the algorithm maintains a commendable degree of stability even as the standard deviation (SD) of noise increases. Specifically, for non-mandatory activities, Acc @3 remains above 80%, indicating a resilient performance against the perturbations caused by noise. With a 5-meter standard deviation in noise, the accuracy at Acc @3 is recorded at 88.4%, which diminishes to 84.2% when the noise level is increased to 20 meters. This reduction, while notable, is not drastic, supporting the assertion that the model upholds a stable predictive capability in the face of increasing noise levels. Particularly with respect to GPS trajectory datasets, the model demonstrates robust inference capabilities at a 5-meter noise accuracy. These findings not only validate the stability of the model but also highlight its potential for effective deployment in real-world scenarios where varying degrees of GPS accuracy are a common occurrence.
TABLE III: Performance metrics for non-mandatory activity inference under different levels of noise
@@
As shown in Table IV, the results indicate that mandatory activities (categories 1-3) exhibit consistent performance due to the deterministic nature of their patterns. On the other hand, activities characterized by more complex and variable patterns present greater challenges for prediction. Despite this, categories 4, 5, 7, 8, 12, and 13 demonstrate commendable accuracy. Notably, categories such as 5 (purchasing goods), 7 (dining out), 12 (availing services), and 13 (religious activities) show robust performance, which can be attributed to the ample instances available within the test dataset.
Conversely, the categories of 11 (visiting friends or relatives), 14 (miscellaneous activities), and 15 (transporting others) prove to be more elusive in the context of stay point detection. These activities are less associated with consistent POIs, leading to relatively poorer model performance. This suggests that the rule-based algorithms are effective for activities with clear-cut patterns.
TABLE IV: Performance metrics for activity inference under 5m noise for different activity types
@@ V Conclusion and Future Work
Our study has introduced a novel LLM-based framework for POI classification and activity inference, demonstrating notable improvements over traditional learning-based methods in handling incomplete and varied datasets like those from OpenStreetMap. Our approach, by leveraging the robust inferential capabilities of large language models, has shown exceptional performance with high accuracy and F-1 scores in real-world scenarios across diverse geographic locations.
Looking ahead, we aim to enhance the accuracy and robustness of POI classification methods with more sophisticated prompt engineering or the development of a fine-tuned LLM designed specifically for the POI classification task.

### v2

@@ IV-B2 Activity Inference
To assess the efficacy of our proposed activity inference model, we curate a test trajectory dataset by extracting activities from the NHTS California Add-on dataset [23], focusing on LA County. This dataset provides detailed daily travel trajectories, with each trajectory consisting of activities and a rough location at the zonal level. However, due to the lack of POIs associated with these activities, we utilize LLM for preliminary POI selection from the rough location zone, followed by manual curation to rectify any erroneous assignments. It is important to note the limitations of GPS-enabled smartphones, which typically provide accurate location data within a 4.9-meter radius under ideal conditions [24]. However, accuracy diminishes in urban environments near structures like buildings, bridges, and dense foliage. To evaluate the model’s performance and robustness against real-world conditions, we introduce noise at three levels: 5 meters, 10 meters, and 20 meters. The noise, simulated using a Gaussian kernel, is added to the latitude and longitude coordinates of the POIs. Subsequent to this augmentation, our test dataset comprises 362 individuals and 2,007 activities, consisting of 1,724 mandatory activities and 283 non-mandatory activities.
@@ IV-C1 POI Classification
As shown in Table II, our model achieves a 93.4% accuracy for the Egypt dataset and 91.4% accuracy for the LA county dataset, with F-1 scores of 96.10% and 92.9% respectively using the gpt4 model. This demonstrates its superiority and robustness compared with previous approaches.
@@ IV-C2 Activity Inference
Table III provides examples of inference and demonstrate the probabilities of activity types and possible visiting POIs. In the example of clear inference, there are two possible POIs near the stay point in the trajectory data. By comprehensively considering the probability of categories and the time of the activity, the model is able to choose ”Buy goods” as the possible activity category, which is consistent with the ground truth. Moreover, the probability of this correct category is significantly higher than in other cases. In contrast, in the example of ambiguous inference, the stay point is located in a commercial area mixed with many shopping and dining POIs. In such a scenario, the model finds ”Buy goods” and ”Buy meals” to be very close, making the situation confusing. The predicted probabilities of ”Buy goods” and ”Buy meals” are very close, leading to a prediction error. The model, even in ambiguous situations, can still make reasonable guesses, although in the case shown in the figure, the second choice is the most accurate inference. Overall, the model’s inference can have strong distinguishability.
Figure 5: Examples of semantic trajectories post-data mining: (a) Activity-annotated stay points in LA. (b) Activity timelines.
TABLE III: Conditional Probability of POI-Activity Combination for a Stay Point*
An example of Clear Inference Activity for POI $A_{1}$ $A_{2}$ $A_{3}$ POI 1 (3.2m)** 0.424 (Buy goods) 0.167 (Buy services) 0.055 (Something else) POI 2 (4.8m) 0.214 (Buy meals) 0.090 (General errands) 0.047 (Something else) An example of Ambiguous Inference Activity for POI $A_{1}$ $A_{2}$ $A_{3}$ POI 1 (3.3m) 0.170 (Buy goods) 0.084 (Buy services) 0.029 (Something else) POI 2 (3.4m) 0.169 (Buy meals) 0.040 (Buy goods) 0.012 (Recreation) POI 3 (3.4m) 0.182 (Buy goods) 0.052 (Buy services) 0.024 (Recreation) POI 4 (4.2m) 0.208 (Buy meals) 0.024 (Recreation) 0 (Buy goods)
* The activity inference algorithm identifies the most likely visited POI within a specified range, accounting for noise and multiple POIs, to determine the final activity type.
** Distance between the POI and the stay point.
According to Table IV, which outlines the performance metrics for activity inference under varying levels of noise, the algorithm maintains a commendable degree of stability even as the standard deviation (SD) of noise increases. Specifically, for non-mandatory activities, Acc @3 remains above 80%, indicating a resilient performance against the perturbations caused by noise. With a 5-meter standard deviation in noise, the accuracy at Acc @3 is recorded at 88.4%, which diminishes to 84.2% when the noise level is increased to 20 meters. This reduction, while notable, is not drastic, supporting the assertion that the model upholds a stable predictive capability in the face of increasing noise levels. Particularly with respect to GPS trajectory datasets, the model demonstrates robust inference capabilities at a 5-meter noise accuracy. These findings not only validate the stability of the model but also highlight its potential for effective deployment in real-world scenarios where varying degrees of GPS accuracy are a common occurrence.
TABLE IV: Performance metrics for non-mandatory activity inference under different levels of noise
@@
As shown in Table V, the results indicate that mandatory activities (categories 1-3) exhibit consistent performance due to the deterministic nature of their patterns. On the other hand, activities characterized by more complex and variable patterns present greater challenges for prediction. Despite this, categories 4, 5, 7, 8, 12, and 13 demonstrate commendable accuracy. Notably, categories such as 5 (Buy goods), 12 (Health care), and 13 (Religious activities) show robust performance, which can be attributed to the ample instances available within the test dataset.
Conversely, activities like 11 (Visit friends), 14 (Something else), and 15 (Drop off/Pick up) prove to be more elusive in the context of stay point detection. These activities are less associated with consistent POIs and often fall into the ambiguous situation as shown in Table III, leading to relatively poorer performance. This suggests that a rule-based algorithm is effective for activities with clear-cut patterns.
TABLE V: Performance metrics for activity inference under 5m noise for different activity types
@@ V Conclusion and Future Work
Our study has introduced a novel LLM-based framework for POI classification and activity inference, demonstrating notable improvements over traditional learning-based methods in handling incomplete and varied datasets like those from OpenStreetMap. Our approach, by leveraging the robust inferential capabilities of large language models, has shown exceptional performance with high accuracy and F-1 scores in real-world scenarios across diverse geographic locations. Looking ahead, we aim to enhance the accuracy and robustness of POI classification methods with more sophisticated prompt engineering or the development of a fine-tuned LLM designed specifically for the POI classification task.